### v2

# Spectral Flow for the Riemann zeros

Abstract
Recently, with Mussardo we defined a quantum mechanical problem of a single particle scattering with impurities wherein the quantized energy levels $E_{n}(\sigma)$ are exactly equal to the zeros of the Riemann $\zeta(s)$ where $\sigma=\Re(s)$ in the limit $\sigma\to\tfrac{1}{2}$. The S-matrix is based on the Euler product and is unitary by construction, thus the underlying hamiltonian is hermitian and all eigenvalues must be real. Motivated by the Hilbert-Pólya idea we study the spectral flows for $\{E_{n}(\sigma)\}$. This leads to a simple criterion for the validity of the Riemann Hypothesis. The spectral flow arguments are simple enough that we present analogous results for the Generalized and Grand Riemann Hypotheses. We also illustrate our results for a counter example where the Riemann Hypothesis is violated since there is no underlying unitary S-matrix due to the lack of an Euler product.
I Introduction
There have been few approaches to the Riemann Hypothesis (RH) based on ideas from physics. Although Riemann’s zeta function $\zeta(s)$ is ubiquitous in Statistical Mechanics, and in Feynman diagram perturbation theory, in this context this usually only involves $\zeta(s)$ with $s$ equal to an integer, however one needs a physical problem which involves $s$ in the entire complex plane. For a review of some approaches based on physics up to the year 2011 see for instance Schumayer . The most prominent physical proposal of the past is probably the Hilbert-Pólya idea which turns the problem of establishing the validity of the Riemann Hypothesis into the existence of a single particle quantum hermitian Hamiltonian whose bound state eigenvalues are equal to the ordinates of zeros on the critical line $\Re(s)=\tfrac{1}{2}$ BerryK ; BerryK2 ; Sierra ; Sredincki ; Bender . The idea is that a non-hermitian hamiltonian would imply complex zeros off the critical line. A primary motivation for this work was the Montgomery-Odlyzko conjecture that the eigenvalue statistics satisfies random matrix GUE statistics for hermitian hamiltonians Montgomery ; Odlyzko . It is fair to say that such a hypothetical hamiltonian has remained elusive, and no specific model has been proposed from which one can calculate the actual zeros.
A different approach, based on statistical physics and random walks, has recently been pursued LMDirichlet ; MLRW ; FL2 . There the validity of the RH is related to random walks over the primes. In this approach one can only prove that the RH is true with probability one, since it relies on the pseudo-randomness of the primes which is difficult to prove. One nice feature of this approach is that the same reasoning applies to all the $L$-functions based on non-principal Dirichlet characters, include those based on cusp modular forms FL2 . This paper does not rely on this work, but is rather focussed in on the Hilbert-Pólya idea for the model in LecMussDefect .
In our recent work with Mussardo LecMussDefect , we proposed a quantum mechanical scattering problem where the quantized energies on a circle correspond to the Riemann zeros in a certain limit. The model consists of a single particle moving on a circle with impurities. It is defined by a real dispersion relation, namely the relation between the momentum $p$ and energy $E$ for a single particle, and the S-matrix for the scattering with the impurities. The model is thus essentially defined by its S-matrix. The S-matrix is unitary thus the model has an implicit underlying hermitian hamiltonian. The main difference with the original version the Hilbert-Pólya idea is that for the latter the quantized energies were imagined to be bound states rather than scattering states. For our model there is a quantization condition for the energies of the system which comes from a Bethe Ansatz equation, and one can thereby readily compute the Riemann zeros accurately from the model, in contrast to some other proposals based on Hilbert-Pólya. The present article continues to study this model, which we will refer to as the LM model.
Let $\rho_{n}=\tfrac{1}{2}+it_{n}$ denote a zero of Riemann’s zeta function $\zeta(s)$ on the upper critical line $t_{n}>0$. In the LM model, quantized energies $E_{n}(\sigma)$ are functions of the abscissa $\sigma$ of the complex variable $s=\sigma+it$. These $E_{n}(\sigma)$ are eigenvalues of a quantum mechanical problem and don’t correspond to zeros of a complex analytic function, except in the limit:
$t_{n}=\lim_{\sigma\to\tfrac{1}{2}^{+}}E_{n}(\sigma).$ (1)
It is thus natural to reconsider the Hilbert-Pólya idea for the LM model, and this is the main focus of this article. The advantage of the LM model is the dependence of the quantized energies on $\sigma$. One is thus led to consider the spectral flow, i.e. how $E_{n}(\sigma)$ depends on $\sigma$. For the LM model, all eigenvalues $E_{n}(\sigma)$ should be real since the underlying hamiltonian is hermitian, and as we will see this will lead a simple criterion for the Riemann Hypothesis (RH) to be true.
Figure 1: A generic situation for eigenvalues of non-hermitian matrices: two real eigenvalues coalesce to form a pair of complex conjugates.
Relatively recently there has been a lot of interest in non-hermitian quantum mechanics in connection with applications to open quantum systems and their Lindbladian dynamics, and some of the ideas of the present work were influenced by this literature. Random matrices for non-hermitian hamiltonians were classified in BL38 of which there are 38 universality classes.bbbThese are now commonly referred to as the Bernard-LeClair classes. Generically, if the $\{E_{n}(\sigma)\}$ are related to a non-hermitian hamiltonian, then a pair of real eigenvalues can coalesce to become a pair of non-real eigenvalues which are complex conjugates. See Figure 1. Although the literature by now is quite large, let us just mention a few works that are perhaps more relevant to this paper Ryu ; Ribiero ; Verbaarschot . In the latter Verbaarschot , at least $14$ of the $38$ classes were realized, and one can find plots for the SYK model SYK1 ; SYK2 exhibiting the generic features displayed in Figure 1. These works are just a motivation and the present work does not rely on this work in detail.cccWe should mention that there are some superficial similarities between the SYK and LM models, in that the interactions are random. In the LM model the randomness stems from the prime numbers that define the S-matrix. However below we will study a Dirichlet $L$-function which corresponds to an underlying non-hermitian hamiltonian and where the RH is thereby false.
We present our work as follows. In the next section we review the LM model and it’s Bethe-Ansatz equations for $\{E_{n}(\sigma)\}$. In Section III we derive the spectral flow equations. This leads to a simple conditions for there to be no complex zeros off the line, which we refer to as Propositions. We present compelling numerical evidence for these propositions. As a test of these ideas it is important to have a counter example which shows how the RH can fail, and this is presented in Section IV. In Section V we show how our propositions are consistent with known results in analytic number theory if one assumes the RH. Since the spectral flow argument so readily leads to equivalences to the RH, in Section VI we present extensions of our propositions to the two infinite classes of $L$-functions: those based on Dirichlet arithmetic characters and those based on (cusp) modular forms. Results from analytic number theory we need can all be found in the elementary books DavenportBook ; Apostol .
II The model and it’s Bethe Ansatz equation
In this section we review the LM model, omitting some details that can be found in LecMussDefect . Consider a single particle of momentum $p$ moving on a circle of circumference $R$. We need to specify a dispersion relation $p(E)$ where $E$ is the energy of a single particle. For now, we chose this dispersion relation to be
$p(E)=E\,\log\left(\dfrac{E}{2\pi e}\right).$ (2)
In the infinite interval for $E>E_{\rm gap}\equiv 2\pi e\approx 17.08$, the above dispersion relation can be inverted as
$E(p)=\dfrac{p}{W(p/2\pi e)}$ (3)
where $W$ is the principal branch of the Lambert $W$ function. Since $E(p)$ must be greater than $2\pi e$, this dispersion relation has an intrinsic energy gap $E_{\rm gap}$. The gap is obvious from the small $p$ expansion,
$\dfrac{E}{2\pi e}=1+\dfrac{p}{2\pi e}-\dfrac{p^{2}}{8\pi^{2}e^{2}}+\dfrac{p^{3}}{12\pi^{3}e^{3}}+{\cal O}(p^{4}).$ (4)
This shows that above the gap $E(p)$ can be interpreted as corrections to a relativistic dispersion. For large $p$, $E(p)\approx p/\log p$, which would be relativistic if it weren’t for the $1/\log p$. This is not a standard dispersion relation, however many uncommon dispersion relations arise in condensed matter physics. For our purposes, (4) implies that the hamiltonian is hermitian since $p$ is a hermitian operator: in position space it is the usual $p=-i\hbar\partial_{x}$.
Consider $N$ stationary impurities spread out on the circle, with no particular location, except that they are separated, and label them $j=1,2,\ldots N$, as illustrated in Figure 2. We assume there is no reflection, namely the scattering is purely transmitting. There are many known examples of purely transmitting relativistic theories DMS ; KonikLeClair ; Corrigan , in fact infinitely many that are integrable, and there are also non-relativistic examples of reflectionless potentials NonRelativistic0 ; NonRelativistic . To each impurity labeled $j$ we associate a transmission S-matrix $S_{j}(p)$, which by unitarity, is a phase
$S_{j}(p)=e^{i\phi_{j}(p)}\,\,\,.$ (5)
Due to the purely transmitting property, the scattering matrix for 2 impurities $j,j^{\prime}$ is simply $S_{j}(p)S_{j^{\prime}}(p)$, and so forth.
Figure 2: Impurities on a circle labeled $j=1,2,...,N$. Different colors denote different scattering phases.
As the particle moves around the circle, it scatters through each impurity and, coming back to its original position, the matching requirement for its wavefunction leads to the quantization condition of its momentum $p$ expressed by
$e^{ipR}\,\prod_{j=1}^{N}S_{j}(p)=\pm 1\,\,\,,$ (6)
where $+1,-1$ corresponds to bosons, fermions respectively (see, for instance MussardoBook and references therein). For our purposes it is necessary to take the particles to be fermions, and we end up with the Bethe-ansatz equation:
$p_{n}R+\sum_{j=1}^{N}\phi_{j}(p_{n})=2\pi\,(n-\tfrac{1}{2})\,,$ (7)
for some integer $n$. Then the quantized energies of the system are $E_{n}=E(p_{n})$, where $p_{n}$ is the solution of the Bethe-ansatz equation relative to the integer $n$. The fact that we need to treat the particles as fermions is perhaps related to the known level-repulsion of energy levels $E_{n}$.
Thus far the model is quite general. For our purposes we define the S-matrices as follows. We assume that the transmission S-matrices are more easily expressed in terms of the energy $E$ rather than the momentum $p$. To each impurity $j=1,2,\ldots$ we associate the $j$-th prime number $\mathfrak{p}_{j}$$\{\mathfrak{p}_{1},\mathfrak{p}_{2},\ldots\}=\{2,3,5,\ldots\}$, and a constant real phase angle $\varphi_{j}$:
$S_{j}(E)=\dfrac{\mathfrak{p}_{j}^{\sigma}-e^{i(E\log\mathfrak{p}_{j}-\varphi_{j})}}{\mathfrak{p}_{j}^{\sigma}-e^{-i(E\log\mathfrak{p}_{j}-\varphi_{j})}},$ (8)
where $\sigma$ is a free parameter which is a positive real number which we will eventually limit to $\sigma>\tfrac{1}{2}$. The phases $\varphi_{j}$ are relevant for more general $L$-functions based on Dirichlet characters. See below for an example based on a mod 5 character. For the Riemann zeta case, we chose $\varphi_{j}=0$. This implies that the scattering phases are
$\phi_{j}(E)=-2\,\Im\log\left(1-\dfrac{e^{-iE\log\mathfrak{p}_{j}}}{\mathfrak{p}_{j}^{\sigma}}\right)\,.$ (9)
Without loss of generality, we set $R=1$, since it can be recovered by scaling the momenta $p$. The Bethe equations (7) read
$\dfrac{E_{n}}{2}\log\left(\dfrac{E_{n}}{2\pi e}\right)-\sum_{j=1}^{N}\Im\log\left(1-\dfrac{e^{-iE_{n}\log\mathfrak{p}_{j}}}{\mathfrak{p}_{j}^{\sigma}}\right)=(n-\tfrac{3}{2})\pi$ (10)
where we have shifted for later convenience $n$ by $-1$.dddThis can be traced to the existence of a pole in $\zeta(s)$ at $s=1$. The solutions $E_{n}$’s obviously depend on $\sigma$, which we denote as $E_{n}(\sigma)$ unless otherwise indicated. Above, the prime numbers $\mathfrak{p}$ could be replaced a real number greater than one, which could be random, and this still defines a physical scattering problem.
Let us now explain how the eigenvalues $E_{n}(\sigma)$ are related to the Riemann zeros. Let us introduce the complex variable $s=\sigma+iE$. If $\sigma>1$, we can use the Euler product for the Riemann zeta function
$\zeta(s)\,=\,\sum_{n=1}^{\infty}\dfrac{1}{n^{s}}\,=\,\prod_{j=1}^{\infty}\left(1-\dfrac{1}{\mathfrak{p}_{j}^{s}}\right)^{-1}~{}~{}~{}~{}~{}(\Re(s)>1).$ (11)
to sum over scattering phases in the thermodynamic limit $N\to\infty$. The equation becomes
$\dfrac{E_{n}}{2}\log\left(\dfrac{E_{n}}{2\pi e}\right)+{\rm arg}\,\zeta(\sigma+iE_{n})=(n-\tfrac{3}{2})\pi.$ (12)
For $\sigma>1$ the thermodynamic limit is well defined since the Euler product converges. As is common in quantum mechanical problems, for $\sigma<1$ we define the sum over scattering phases via analytic continuation, and this thermodynamic limit through analytic continuation will be henceforth assumed. It is important that ${\rm arg}$ in the above equation is not the principle branch ${\rm Arg}$; ${\rm arg}$ is a sum of the ${\rm Arg}$’s in (9), but they can accumulate. In fact it is known that the above ${\rm arg}$ is unbounded, being roughly $\log E$. If one neglects the ${\rm arg}\,\zeta$ term, then for large $n$ the solution to the equation (12) above is approximately $E_{n}\approx\widetilde{E}_{n}$ where
$\widetilde{E}_{n}\approx\dfrac{2\pi(n-\tfrac{11}{8})}{W\left((n-\tfrac{11}{8})/e\right)}$ (13)
where $W$ is the principle branch of the Lambert function. The $\widetilde{E}_{n}$ should be viewed as the quantized energies of the free theory with the S-matrix equal to 1.
The above equation (12) was first proposed in Electrostatic , and is an excellent approximation to the Riemann zeros on the critical line as $\sigma\to\tfrac{1}{2}^{+}$. See (20) below. The reason for this is quite simple as we now explain. Following standard conventions in analytic number theory, let us now define a complex variable $s=\sigma+it$ where, based on the notation above, $t=E$. Henceforth we will use $E,t$ interchangeably if there is no cause for confusion. We consider zeros on the critical line, which are known to be infinite in number. Denote the $n$-th zero on the upper critical line as
$\rho_{n}=\tfrac{1}{2}+it_{n},~{}~{}~{}~{}~{}n=1,2,3,\ldots,~{}~{}~{}\zeta(\rho_{n})=0$ (14)
where $t_{1}=14.1347..$ is the first zero, and so forth. Define a completed $\zeta$ function which satisfies the functional equation as follows:
$\chi(s)=\pi^{-s/2}\Gamma(s/2)\zeta(s),~{}~{}~{}~{}~{}\chi(1-s)=\chi(s).$ (15)
Furthermore, define its argument $\theta$:
$\theta(\sigma,t)={\rm arg}\,\chi(\sigma+it),~{}~{}~{}~{}~{}~{}\theta(\sigma,t)=\vartheta(\sigma,t)+{\rm arg}\,\zeta(\sigma+it),~{}~{}~{}~{}~{}~{}~{}~{}\vartheta(\sigma,t)\equiv{\rm arg}\,\Gamma\left(\tfrac{1}{2}(\sigma+it)\right)-\tfrac{t}{2}\log\pi.$ (16)
On the critical line $\vartheta(\tfrac{1}{2},t)$ is smooth and commonly referred to as the Riemann-Siegel $\vartheta$ function. Below, if it is implicit that we are on the critical line we will simply write $\vartheta(t)\equiv\vartheta(\tfrac{1}{2},t)$. As argued in LecMussDefect , since $\chi(s)$ is real on the critical line, it must jump by $\pi$ at each simple zero. This implies that the exact Riemann zeros on the line
$t_{n}=\lim_{\sigma\to\tfrac{1}{2}^{+}}\,E_{n}(\sigma),$ (17)
satisfy the equation FrancaLeClair ; LecMussDefect :
$\lim_{\delta\to 0^{+}}\theta(\tfrac{1}{2}+\delta,t_{n})=(n-\tfrac{3}{2})\pi,~{}~{}~{}~{}\Longrightarrow~{}~{}~{}\vartheta(t_{n})+\lim_{\delta\to 0^{+}}{\rm arg}\,\zeta(\tfrac{1}{2}+\delta+it_{n})=(n-\tfrac{3}{2})\pi.$ (18)
Let us now return to the equation (12). One easily sees that the $\vartheta(t_{n})$ term in (18) amounts to small corrections to our dispersion relation (2):
$p(E)\approx 2\vartheta(E)=E\,\log\left(\tfrac{E}{2\pi e}\right)-\dfrac{\pi}{4}+\dfrac{1}{24E}+O(1/E^{3}).$ (19)
As for $p(E)$ in eq. (2), $\vartheta(E)$ is monotonic and positive in the infinite interval $(E_{*},\infty)$, where $E_{*}\simeq 17.8456$ is a minor deformation of the gap $E_{\rm gap}=2\pi e$, and in this interval it is invertible and therefore asymptotically eq.(3) is valid. Thus (12) should be viewed as a large $t$ asymptotic version of the exact equation (18). However for reasons already mentioned, it is quite accurate for relatively low $t$, and (12) provides excellent results for even the first few zeros. For instance from (12) with $\delta=10^{-6}$ we obtain
$\lim_{\sigma\to\tfrac{1}{2}^{+}}\{E_{1}(\sigma),E_{2}(\sigma),\ldots,E_{5}(\sigma)\}=\{14.134724,~{}21.022039,~{}25.010857,~{}30.424875,~{}32.935061\},$ (20)
which are identical to the true Riemann zeros to 8 digits. We emphasize that the exact equation (18) yields zeros to arbitrary precision by systematically reducing $\delta$, as illustrated in FrancaLeClair . Henceforth, we will work mainly with the asymptotic equation (12) since the $1/E$ corrections in (19) are either negligible for high enough $E_{n}$ or easily incorporated without affecting the main results; such corrections just make the formulas more complicated in an irrelevant manner for our purposes. (See for instance equation (43) where we display such corrections in that case.) Thus, below the symbols $\lesssim$ and $\simeq$ refer merely to neglecting these ${\cal O}(1/E)$ corrections. The justification for this will be clear from the numerical results below.
The above considerations lead to our first Proposition:
Proposition 1.
Since the S-matrix for the LM model, which is based on the Euler product formula, is unitary, and the non-interacting part of the hamiltonian is hermitian based on (4), all eigen-energies $E_{n}(\sigma)$ are real in the thermodynamic limit with $\sigma>\tfrac{1}{2}$.
The LM model is mainly defined by it’s unitary S-matrix based on the Euler product. Although we didn’t present a formula for the implicit underlying hamiltonian, there are many examples of such theories defined by their S-matrix, namely where the theory is defined by its S-matrix and an explicit hamiltonian is unknown. Henceforth we assume the theory is well-defined by its dispersion relation and it’s unitary S-matrix.
In LecMussDefect we argued that the RH is true if $\chi(s)$ is a regular alternating function, however after some attempts this seems difficult to prove. The present article will present much more succinct criteria for validity of the RH.
Let us mention a relevant remark since it is one of the primary motivations for this article. One has the following theorem FrancaLeClair :
$\bullet$   If for every integer $n>0$ there is a unique solution of the equation (18) for $E_{n}(\sigma)$ in the limit $\sigma\to\tfrac{1}{2}^{+}$, then the Riemann Hypothesis is true and all zeros are simple.
The reason this statement is correct is quite simple: If there is a solution for every $n$, then the zeros along the critical line are enumerated and their counting saturates the known formula for the total number of zeros in the critical strip up height $t<T$, commonly referred to as $N(T)=\vartheta(T)/\pi+{\rm arg}\,\zeta(\tfrac{1}{2}+it)/\pi+1$. Clearly any zeros off the critical line will spoil this. In Section IV we show how this is violated for a counterexample.
III Spectral flow and a simple criterion for the validity of the Riemann Hypothesis
III.1 The spectral flow equation
As in the last section, we will use $E$ and $t$ interchangeably. We will mainly work with the asymptotic form (12) rather than the exact (18) since order $1/E_{n}$ corrections are easily neglected as explained above. This will be implicit in some formulas below, where by “large” $t$ we mean $t\gtrsim 20$.
For $\sigma>1$ there is a unique solution $E_{n}(\sigma)$ to equations (18) and (12) for each $n$ since the ${\rm arg}\,\zeta$ term is well defined through the Euler product and the left hand side is monotonic. In this section we are concerned with the dependence of $E_{n}(\sigma)$ on $\sigma$ as one moves into the critical strip $\tfrac{1}{2}<\sigma<1$ starting from $\sigma>1$ where the $E_{n}(\sigma)$ are well-defined. In Figure 3 we show this dependence on $\sigma$ for the lowest five zeros. One sees that the $E_{n}$ smoothly deform to the actual zeros $t_{n}$ as $\sigma\to\tfrac{1}{2}$ without a great deal of variation nor drama. Compare with Figure 10 below, where as we will see the RH is violated. This implies the zeros $t_{n}$ can be roughly estimated by $E_{n}(\sigma)$ with $\sigma>1$ where the Euler product converges.
Figure 3: A plot of $E_{n}(\sigma)$ for $n=1,2,\ldots,5$ as a function of $\sigma$ to the right of the critical line.
As motivated in the Introduction, if there were a zero off the critical line, one expects a pair of real eigenvalues to coalesce into a pair of complex eigenvalues. Inspection of Figure 1 indicates that $dE_{n}/d\sigma$ diverges at this point. Differentiating (18) with respect to $\sigma$ one finds
$\dfrac{dE_{n}(\sigma)}{d\sigma}=-\dfrac{\Im(\Upsilon(s))}{\Re(\Upsilon(s))+\vartheta^{\prime}(E_{n}(\sigma))},~{}~{}~{}~{}~{}s=\sigma+iE_{n}(\sigma)$ (21)
where $\vartheta^{\prime}(E)=\partial_{E}\vartheta(E)$ and we have defined
$\Upsilon(s)=\dfrac{\zeta^{\prime}(s)}{\zeta(s)},~{}~{}~{}~{}~{}~{}\zeta^{\prime}(s)=\partial_{s}\zeta(s).$ (22)
We thus propose that at a zero off the critical line, the denominator in (21) vanishes. This leads us to the proposition:
Proposition 2.
If for a given $\sigma_{*}$ and a region $t_{1}<t<t_{2}$, the following condition is satisfied
$-\Re(\Upsilon(s))<\vartheta^{\prime}(t),~{}~{}~{}~{}~{}s=\sigma_{*}+it,$ (23)
then all $E_{n}(\sigma)$ with $\sigma>\sigma_{*}$ in this region of $t$ are real. For large enough $t\gtrsim 20$, the above is equivalent to
$-\Re(\Upsilon(s))\lesssim\,\dfrac{1}{2}\log\left(\dfrac{t}{2\pi}\right).$ (24)
Clearly this implies that the RH is true if the condition (24) is satisfied for all $\sigma>\tfrac{1}{2}$ and for all $t>0$.
One can also learn something by considering the integer $n$ as a continuous variable. On the critical line, if the $E_{n}(\sigma)$ are real solutions, then they are discrete and $dE_{n}(\sigma)/dn$ should be zero between zeros. Differentiating (12), one finds
$\dfrac{\partial E_{n}(\sigma)}{\partial n}=\dfrac{\pi}{\Re(\Upsilon(s))+\tfrac{1}{2}\log\left(E_{n}(\sigma)/2\pi\right)},~{}~{}~{}~{}~{}~{}s=\sigma+iE_{n}(\sigma).$ (25)
The same denominator as in (21) appears. Now if the RH is true, all $E_{n}(\sigma=\tfrac{1}{2})$ are real and well-defined and $dE_{n}/dn$ must be zero between zeros. This leads to another proposition:
Proposition 3.
For any $t$ not equal to a zero $t_{n}$ on the critical line, one has the equality:eee We repeat that it is implicit that this is for $t$ large enough, namely $\simeq$ in the above equation signifies we are neglecting ${\cal O}(1/t^{2})$ corrections which vanish at large $t$ and don’t affect our Propositions for $t\gtrsim 20$.
$-\Re(\Upsilon(s))\simeq\tfrac{1}{2}\log(t/2\pi),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}s=\tfrac{1}{2}+it.$ (26)
Since this is an equality and not a bound, the $1/2\pi$ inside the $\log$ is essential.
Thus far we have not assumed the RH, and this is of key importance. In fact, Propositions 2 and 3 follow from Proposition 1. As we will explain in Section V, Proposition 2 has already been shown to be equivalent to the RH by Lagarias Lagarias using some rather sophisticated analysis compared to our simple spectral flow argument, and of course by assuming the RH.fffWe discovered this subsequent to the above spectral flow argument upon searching through the equivalents of the RH listed in Borwein . In Section V, we will provide a relatively simple proof of the above Proposition 2 by assuming the RH.
III.2 Numerical evidence
In this section we provide some convincing numerical evidence for the above Propositions, in particular in the forms (24) and (26).
In Figure 4 we consider $\sigma=\tfrac{3}{4}$ for very low $t$. Clearly above the relatively low value $t\sim 10$, the inequality (24) is satisfied. It is interesting to consider the same $\sigma=\tfrac{3}{4}$ for a considerably higher $t\sim 1000$ as shown in Figure 5. Both these plots rule zeros for $\sigma>\tfrac{3}{4}$ in these ranges of $t$, but interestingly it is easier to rule out such zeros for higher $t$.
In Figure 6 we approach the critical line with $\sigma=\tfrac{1}{2}+10^{-4}$. As expected one gets much closer to the bound (24), but it is still satisfied. This pattern repeats itself at larger $t\sim 1000$, as shown in Figure 6. What is interesting is that exactly on the critical line, the zeros are hardly visible and fit equation (24) almost perfectly, supporting Proposition 3.
Figure 4: A plot $-\Re(\Upsilon(s))$ for $s=\tfrac{3}{4}+it$ and $\tfrac{1}{2}\log(t/2\pi)$ as a function of $t$ for $300<t<$ 30. The asymptotic version of our bound (24) is only violated for very low $t\lesssim 10$ which is uninteresting.
Figure 5: A plot $-\Re(\Upsilon(s))$ for $s=\tfrac{3}{4}+it$ and $\tfrac{1}{2}\log(t/2\pi)$ as a function of $t$ for $10051000<t<$ 1005.
Figure 6: A plot $-\Re(\Upsilon(s))$ for $s=\tfrac{1}{2}+\delta+it$ with $\delta=10^{-4}$ and $\tfrac{1}{2}\log(t/2\pi)$ as a function of $t$ for $300<t<$ 30.
Figure 7: A plot $-\Re(\Upsilon(s))$ for $s=\tfrac{1}{2}+\delta+it$ with $\delta=10^{-7}$ and $\tfrac{1}{2}\log(t/2\pi)$ as a function of $t$ for $10051000<t<$ 1005.
Figure 8: A plot $-\Re(\Upsilon(s))$ on the critical line $s=\tfrac{1}{2}+it$ and $\tfrac{1}{2}\log(t/2\pi)$ as a function of $t$ for $10051000<t<$ 1005. The zeros are points on this line which are invisible at this resolution.
III.3 Analytic continuation of the Euler product for $\Upsilon(s)$.
The Euler product has been a crucial ingredient in the above considerations. Compared to $\log\zeta$ and ${\rm arg}\,\zeta$, it is easier to analytically continue $\Upsilon$ into the strip and it is better behaved there. For this subsection let us indicate one analytical continuation of $\Upsilon(s)$ into the strip based on the prime-zeta function. We will not use this result in the sequel, rather presenting it is meant to show that $\Upsilon$, originally defined over primes (27), can be analytically continued in and of itself into the critical strip.
$-\Upsilon(s)=\sum_{\mathfrak{p}}\dfrac{\log\mathfrak{p}}{\mathfrak{p}^{s}-1},~{}~{}~{}~{}~{}\Re(s)>1.$ (27)
Define the prime-zeta function $P(s)$:
$P(s)=\sum_{\mathfrak{p}}\dfrac{1}{\mathfrak{p}^{s}}.$ (28)
Then expanding $1/(\mathfrak{p}^{s}-1)$ one obtains the known formula
$\Upsilon(s)=\sum_{n=1}^{\infty}P^{\prime}(ns),$ (29)
for $\Re(s)>1$. Whereas the sum over primes in (27) does not converge, it is known how to analytically continue $P(s)$ into the critical strip, and the sum over $n$ in (29) converges.
IV Spectral flow for a non-hermitian example which violates the RH
In order to sharpen the above ideas, it is very useful to have an example which has almost all of the properties of $\zeta(s)$ but the RH is violated. There is a well-known such example due to Davenport-Heilbronn DavenportH . It satisfies a functional equation like $\zeta(s)$, it has an infinite number of zeros on the line and also off the line. As we explain, the main difference is that it does not have an Euler product representation. Recall that for $\zeta(s)$ the Euler product was necessary to define a unitary S-matrix. Without the Euler product, one cannot even define the quantum scattering problem as in Section II LecMussDefect . Without a unitary S-matrix, the underlying hamiltonian must be non-hermitian, and this implies that there should be complex eigenvalues which violate the RH. In this section we study the spectral flow for the eigenvalues $E_{n}(\sigma)$ and show that indeed the analog of Proposition 2 is violated.
Let $L(s,X)$ denote the Dirichlet $L$-function based on the non-principal and primitive mod $q=5$ character
$\{X(1),X(2),\ldots,X(5)\}=\{1,i,-i,-1,0\},~{}~{}~{}~{}~{}(q=5).$ (30)
Due to the complete multiplicity of the characters, $X(n)X(m)=X(nm)$, it has both a series representation and an Euler product as in (31):
$L(s,X)=\sum_{n=1}^{\infty}\dfrac{X(n)}{n^{s}}\,=\,\,\prod_{j=1}^{\infty}\left(1-\dfrac{X(\mathfrak{p})}{\mathfrak{p}_{j}^{s}}\right)^{-1}~{}~{}~{}~{}~{}(\Re(s)>1).$ (31)
The DH function ${\cal D}(s)$ is constructed to satisfy a duality functional equation. It is defined by the linear combination
${\cal D}(s)\equiv\tfrac{(1-i\kappa)}{2}\,L(s,X)+\tfrac{(1+i\kappa)}{2}\,L(s,\overline{X})$ (32)
where $\kappa\equiv\tfrac{\sqrt{10-2\sqrt{5}}-2}{\sqrt{5}-1}$. Each $L(s,X)$ function in the definition (32) has the Euler product formula representation in (31), however the linear combination in ${\cal D}(s)$ does not.
For the remainder of this section we use the same notation $\chi(s)$, $\theta(\sigma,t)$, $\vartheta(\sigma,t)$, $t_{n}$, $E_{n}(\sigma)$, and $\Upsilon$ as above for $\zeta$, however they all refer to the DH function. Define it’s completion
$\chi(s)\equiv\left(\tfrac{\pi}{5}\right)^{-s/2}\Gamma\left(\tfrac{1+s}{2}\right)\,{\cal D}(s),$ (33)
which satisfies $\chi(s)=\chi(1-s)$ by design. As for $\zeta(s)$ we consider its argument:
$\theta(\sigma,t)={\rm arg}\,\chi(s)=\vartheta(\sigma,t)+{\rm arg}\,{\cal D}(s);~{}~{}~{}~{}~{}\vartheta(\sigma,t))={\rm arg}\,\Gamma\left(\tfrac{s+1}{2}\right)-\tfrac{t}{2}\log(\tfrac{\pi}{5}).$ (34)
The appropriate Riemann-Siegel $\vartheta$ on the critical line is $\vartheta(t)\equiv\vartheta(\tfrac{1}{2},t)$. The analog of eqn. (18) for the hypothetical $n$-th zero $\rho=\tfrac{1}{2}+it_{n}$ of the DH function on the critical line is simply FrancaLeClair :
$\lim_{\delta\to 0^{+}}\theta(\tfrac{1}{2}+\delta,t_{n})=\left(n-\tfrac{1}{2}\right)\pi.$ (35)
The counting function for the number of zeros up to height $T$ in the entire critical strip as usual follows from the argument principle:
$N(T)=\lim_{\delta\to 0^{+}}\theta(\tfrac{1}{2}+\delta,T)/\pi.$ (36)
As for $\zeta(s)$ we deform the equation to obtain equations for $E_{n}(\sigma)$ where $t_{n}=\lim_{\sigma\to\tfrac{1}{2}^{+}}E_{n}(\sigma)$. (For more details see LecMussDefect .) In the limit of large enough $E_{n}$
$\dfrac{E_{n}(\sigma)}{2}\log\left(\dfrac{q\,E_{n}(\sigma)}{2\pi e}\right)+{\rm arg}\,{\cal D}(\sigma+iE_{n}(\sigma))=(n-\tfrac{1}{2})\pi.$ (37)
The sectral flow equation is of the form (21). Repeating arguments from Section III, the RH for this $L$-function would be true if the following condition is satisfied
$-\Re\left(\Upsilon(s)\right)<\dfrac{1}{2}\log\left(\dfrac{5t}{2\pi}\right),~{}~{}~{}{\rm for}~{}~{}\Re(s)>\tfrac{1}{2},~{}~{}~{}~{}\Upsilon(s)\equiv\dfrac{{\cal D}^{\prime}(s)}{{\cal D}(s)}.$ (38)
For illustration, let us limit ourselves to the range $1000<t<$ 100. One can easily check numerically that (35) correctly gives the first $43$ zeros, which are all on the critical line. However for $n=44$ and $45$ there is no solution to the equation (35) FrancaLeClair ; LecMussDefect . This can be seen from Figure 9, and signifies two zeros off the line which are complex conjugates. They are:
[…]
There are an infinite number of zeros on the line which satisfy (35), and also an infinite number of zeros off the line where there are no solutions to (35). The next $n$ where the latter occurs is $n=63$.
Figure 9: A plot of $N(t)$ in (36) as a function of $t$ in the range of the first zero off the line. We emphasize that the non-zero $\delta$ in (36) is necessary otherwise $N(t)$ would jump discontinuously by $1$ at zeros on the line.
Let us now see how this is consistent with the spectral flow condition (38), namely specifically how it is violated. In Figure 10 we plot $E_{n}(\sigma)$ as a function of $\sigma$ for $n$ in the vicinity of zeros off the line. Near the zero off the line, one observes what was anticipated in the Introduction, namely a pair of eigenvalues $E_{n}$ which are well-defined for $\sigma\gtrsim 0.81$ coalesce into a pair of complex eigenvalues. Closer inspection in Figure 11 suggests a level-crossing. In any case $dE_{n}/d\sigma$ is infinite there. Furthermore, the $E_{n}(\sigma)$ are not well defined as $\sigma\to\tfrac{1}{2}^{+}$, as can also be seen in the Figure 10.
Figure 10: A plot of $E_{n}(\sigma)$ as a function of $\sigma$ for $n=42,43,44,45,46,47$ for the Davenport-Heilbronn function. One sees that $E_{n}(\sigma)$ is ill-defined as $\sigma\to\tfrac{1}{2}$ for $n=44,45$.
[…]
In the range $1000<t<$ 100 there are no zeros off the line with $\sigma>0.9$, thus (38) should be valid, and it is as can been seen from Figure 12. On the other hand, for $\sigma>0.7$ the condition (38) should be violated due to the zero in (39), and indeed it clearly is, as can be seen in Figure 13.
Figure 12: A plot of $-\Re(\Upsilon(\sigma+it)$ verses $\tfrac{1}{2}\log(5t/2\pi)$ for the Davenport-Heilbronn function as a function of $t$ for $\sigma=0.9$.
Figure 13: A plot of $-\Re(\Upsilon(\sigma+it)$ for the Davenport-Heilbronn function verses $\tfrac{1}{2}\log(5t/2\pi)$ as a function of $t$ for $\sigma=0.7$.
V Analytic proof of the Propositions for $\zeta(s)$ assuming the RH
Thus far we have argued for Propositions 2, 3 without assuming the RH, namely it would follow from the hermiticity of the underlying hamiltonian of the LM model due to its unitary S-matrix as stated in Proposition 1. In this section we show that the Propositions of Section III can be derived if one assumes the RH.
Lagarias proved Lagarias that the RH is true if
$\Re\left(\dfrac{\xi^{\prime}(s)}{\xi(s)}\right)>0~{}~{}~{}~{}~{}{\rm for}~{}\Re(s)>\tfrac{1}{2},~{}~{}~{}~{}~{}\xi(s)\equiv s(s-1)\,\chi(s)/2.$ (40)
Using known results found in DavenportBook , it can be shown that this implies Proposition 2. Let us show this directly in a somewhat simpler fashion by working directly with $\zeta(s)$, although this argument must be implicit in Lagarias’ work. By taking a logarithmic derivative of the functional equation $\xi(s)=\xi(1-s)$ and the Hadamard product, and also utilizing the Euler product, one can show DavenportBook
$-\Upsilon(s)=\dfrac{1}{s-1}-B-\log\sqrt{\pi}+\dfrac{1}{2}\dfrac{\Gamma^{\prime}(s/2+1)}{\Gamma(s/2+1)}-\sum_{\rho}\left(\dfrac{1}{s-\rho}+\dfrac{1}{\rho}\right),$ (41)
where
$B=-\gamma_{E}/2-1+\log(2\sqrt{\pi})=-0.0230957...$ (42)
with $\gamma_{E}$ the Euler constant. Above, $\sum_{\rho}$ denotes a sum over all non-trivial zeros inside the critical strip $10<\Re(\sigma)<$ 1, including potential zeros off the critical line. Using the Stirling formula,
$-\Re(\Upsilon(s))=\tfrac{1}{2}\log(t/2\pi)\,+\,\Re\left(\dfrac{1}{s-1}\right)-B+\dfrac{(2+6\sigma+3\sigma^{2})}{12t^{2}}-\Re\,\sum_{\rho}\left(\dfrac{1}{s-\rho}+\dfrac{1}{\rho}\right)+{\cal O}(1/t^{4}).$ (43)
Let us turn to $\sum_{\rho}$. Given a zero at $\rho_{\bullet}=\sigma_{\bullet}+it_{\bullet}$, the functional equation and complex conjugation implies there are a total of 4 zeros at $\rho_{\bullet}=1-\sigma_{\bullet}\pm it_{\bullet}$ and $\rho_{\bullet}=\sigma_{\bullet}\pm it_{\bullet}$. Now $\Re(1/\rho_{\bullet})=\sigma_{\bullet}/|\rho_{\bullet}|^{2}$. Thus in the sum over $\rho_{\bullet}$, zeros off the line $\sigma_{\bullet}\neq\tfrac{1}{2}$ cancel and one is left with a sum over zeros on the critical line:
$\sum_{\rho}\,\Re\left(\dfrac{1}{\rho}\right)=\sum_{n}\dfrac{1}{|\rho_{n}|^{2}},~{}~{}~{}~{}~{}\rho_{n}=\tfrac{1}{2}+it_{n},~{}~{}~{}~{}t_{n}>0$ (44)
The above sum actually converges and cancels the $B$ term in (43). For instance, for the first 200,000 zeros we find
$\sum_{n=1}^{2\cdot 10^{5}}\,\dfrac{1}{|\rho_{n}|^{2}}=0.0230832.$ (45)
For $\tfrac{1}{2}<\sigma<1$, one sees that $\Re(1/(s-1))<0$. Finally one has
$\Re\left(\dfrac{1}{s-\rho_{\bullet}}\right)=\dfrac{\sigma-\sigma_{\bullet}}{(\sigma-\sigma_{\bullet})^{2}+(t-t_{\bullet})^{2}}.$ (46)
If there is a zero off the line with $\sigma_{\bullet}>\tfrac{1}{2}$ then for $\sigma<\sigma_{\bullet}$ the above term is negative and this spoils the bound (24). Assuming the RH, all $\sigma_{\bullet}=\tfrac{1}{2}$ and the above is positive for $\sigma>\tfrac{1}{2}$, thus the bound (24) is satisfied. This proves Proposition 2 in the form (24).
Again assuming the RH, on the critical line the expression (46) vanishes. Neglecting all $1/t^{2}$ terms, including those from $\Re(1/(s-1))$, which is valid for large $t$, one obtains Proposition 3.
VI Generalized and Grand Riemann Hypothesis
The Riemann zeta function is a special case of an infinite class of $L$-functions based on Dirichlet characters; it corresponds to that of the principal character mod $q=1$. It has been conjectured that the RH is true for all such functions, and this is referred to as the Generalized Riemann Hypothesis. These functions enjoy both a functional equation relating $s$ to $1-s$, and an Euler product representation. A completely different class of $L$-functions are built upon the Fourier coefficients of modular forms, and also satisfy a functional equation and Euler product. The conjecture that they also satisfy a RH is referred to as the Grand Riemann Hypothesis. The random walk argument for these class of functions was presented in LMDirichlet ; MLRW ; FL2 . Here we focus on arguments based on the LM model, which are easily generalized to these two infinite classes of $L$-functions.
The above spectral flow equations are easily generalized with little detailed analysis to these classes of functions, thus we present the analogs of the above Propositions to these cases. Because of the existence of the Euler product, one can easily define the impurity scattering problem as we did for $\zeta(s)$ for a finite number $N$ of impurities, define the quantized energies $E_{n}(\sigma)$, and take the thermodynamic limit. One thereby obtains the equations for individual zeros on the critical line $\Re(s)=\tfrac{1}{2}$ proposed in FrancaLeClair with $\sigma\neq\tfrac{1}{2}$. Based on such equations it is straightforward to obtain simple criteria for the RH to hold based on the spectral flow argument of Section III. As above for Riemann $\zeta$, we drop ${\cal O}(1/t^{2})$ corrections to the dispersion relation coming from the analog of Riemann-Siegel $\vartheta$, since they are negligible for large enough $t$, which in fact is not so large, again around $t\gtrsim 10$. Below the statement “for $t$ large enough” refers to this.
VI.1 Generalized Riemann Hypothesis
Let $L(s,X)$ denote the $L$-function based on a primitive Dirichlet character $X$ of modulus $q$. Due to the complete multiplicity of the characters $X(n)X(m)=X(nm)$, it has both a series and a Euler product representation:
$L(s,X)=\sum_{n=1}^{\infty}\dfrac{X(n)}{n^{s}}\,=\,\,\prod_{j=1}^{\infty}\left(1-\dfrac{X(\mathfrak{p})}{\mathfrak{p}_{j}^{s}}\right)^{-1}~{}~{}~{}~{}~{}(\Re(s)>1).$ (47)
The equations in FrancaLeClair for individual zeros on the line depend on various details such as the Gauss sum and other number-theoretic quantities, but for our purposes these are negligible for large enough $t$.gggSee for instance equation (43) in FrancaLeClair . Repeating the arguments of Section III we can propose the following.
Proposition 4.
The Generalize Riemann Hypothesis is true if for $t$ large enough
$-\Re\left(\Upsilon(s)\right)<\dfrac{1}{2}\log\left(\dfrac{qt}{2\pi}\right),~{}~{}~{}~{}\Upsilon(s)\equiv\dfrac{L^{\prime}(s,X)}{L(s,X)}$ (48)
for $\Re(s)>\tfrac{1}{2}$.
VI.2 Grand Riemann Hypothesis
For a detailed exposition of the mathematical background for the $L$-functions in this section, we refer to Apostol’s book Apostol . Let $f(\tau)$ denote a cusp form of weight $k$. By definition under an $SL(2,\mathbb{Z})$ tranformation:
$f\left(\dfrac{a\tau+b}{c\tau+d}\right)=(c\tau+d)^{k}\,f(\tau).$ (49)
One $SL(2,\mathbb{Z})$ tranformation implies the periodicity $\tau\to\tau+1$, thus it has a Fourier series
$f(\tau)=\sum_{n=0}^{\infty}a_{f}(n)\,q^{n},~{}~{}~{}~{}~{}q=e^{2\pi i\tau}.$ (50)
From these Fourier coefficients one defines the $L$-function
$L_{f}(s)=\sum_{n=1}^{\infty}\dfrac{a_{f}(n)}{n^{s}}.$ (51)
It is well-known that it satisfies a functional equation and Euler product formula DavenportBook , and the critical line is $\Re(s)=k/2$. The Euler product is slightly more complicated since the coefficients $a_{f}(n)$ are not completely multiplicative, but this does not affect our reasoning. In FrancaLeClair the analog of exact equations satisfied by individual zeros were presented, and this is all we need for our purposes:
$t_{n}\log\left(\dfrac{t_{n}}{2\pi e}\right)+\lim_{\delta\to 0^{+}}{\rm arg}\,L_{f}(\tfrac{k}{2}+\delta+it_{n})=\left(n-\dfrac{k+(-1)^{k/2}}{4}\right)\pi.$ (52)
This leads to the Proposition:
Proposition 5.
The Grand Riemann Hypothesis is true if for $t$ large enough
$-\Re\left(\Upsilon(s)\right)<\log\left(\dfrac{t}{2\pi}\right),~{}~{}~{}~{}~{}\Upsilon(s)\equiv\dfrac{L^{\prime}_{f}(s)}{L_{f}(s)}$ (53)
for $\Re(s)>k/2$.
Note there is no $k$ dependence of the above bound, and it differs from the $\zeta(s)$ case only in the overall coefficient of the $\log$ which is $1$ rather than $\tfrac{1}{2}$. The statistical mechanical approach involving random walks was described in FL2 .
VII Conclusions
We proposed that the quantized energies $\{E_{n}(\sigma)\}$ for the LM model are real due to the fact that its S-matrix is unitary based on the Euler product and the hamiltonian for the free part, i.e. the dispersion relation, is hermitian (Proposition 1.) Without assuming the RH, we studied the spectral flow, namely how the $E_{n}(\sigma)$ depend on $\sigma$ and derived a simple condition for all eigenvalues to be real, stated as Propositions 2 and 3 above. We showed how these Propositions are equivalent to the validity of the Riemann Hypothesis. We extended this reasoning to the two infinite classes of functions for the Generalized and Grand Riemann Hypotheses. We also presented spectral flow arguments for a counterexample where the RH is violated, and attributed this to the lack of an Euler product which implies a non-unitary S-matrix. This work provides ample evidence that for the RH to be true, one needs both the functional equation and the Euler product.
Based on this article, the validity of the RH comes down Proposition 1, namely that the hamiltonian for the LM model and its generalizations has real eigenvalues $E_{n}(\sigma)$. We argued that this is valid due to the unitarity of the S-matrix and the reality and invertibility of the dispersion relation. In fact, it is clear from (4) that the free hamiltonian for our dispersion relation is a hermitian operator. We thus cannot see any obstructions to 1 at present. We point out that the recent work Negro seems relevant to the study of the free non-interacting part of the hamiltonian if the dispersion relation is based on the exact Riemann-Siegel $\vartheta$ function.
An appealing feature of the above Propositions is that they all take a universal form, namely that the appropriate quantity, $-\Re(\Upsilon(s))$, is bounded by simple variations of $\log(t/2\pi)$ for both the Generalized and Grand Riemann Hypotheses, and this suggests a simple, universal strategy towards understanding all of them. This is in contrast to most ideas based on Hilbert-Pólya which are idiosyncratic to only $\zeta(s)$ itself and do not indicate how to proceed beyond to these Generalized and Grand Riemann Hypotheses.
There are many interesting avenues for further exploration. The $\{E_{n}(\sigma)\}$ are well defined for all $\sigma>\tfrac{1}{2}$, and it would be interesting to understand if they still satisfy GUE random matrix statistics as is the case for $\sigma=\tfrac{1}{2}$. We suspect that they do, since the S-matrix that defines the model is based on the primes which are pseudo-random. For the $L$-functions that violate the RH, such as the example in Section IV, it would be interesting to study if the complex eigen-energies satisfy the statistics of random non-hermitian hamiltonians classified in BL38 .
The LM model is more general than the specific realization we studied here where the S-matrix was defined using the prime numbers. However one can replace the primes with other real numbers and still have a well-defined scattering problem with impurities LecMussDefect . If the impurities are random phases then this defines an interesting problem in Anderson localization. For the version of the LM model in this paper, the randomness of the impurities stems from the pseudo-randomness of the primes.
VIII Acknowledgements
We especially thank Giuseppe Mussardo for our collaboration in LecMussDefect which led to the considerations of this paper. We also thank the Simons Center for Geometry and Physics and the organizers for their invitation to their recent program in May 2024 on far from equilibrium physics which stimulated some of these ideas, in particular conversations with J. Verbaarschot.